The distance a body falls caries directly as the square of the time that it falls. If it falls 972 meters in 18 seconds, how far will it fall in 2 seconds?
Answer by Alan3354(69443)

The distance a body falls caries directly as the square of the time that it falls. If it falls 972 meters in 18 seconds, how far will it fall in 2 seconds?
------------
972 = k*18^2 = 324k
k = 3
=============
s = kt^2 = 3*2^2
s = 12 meters in 2 seconds